OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#8717 — Collection 13
Scheduled Maintenance Report for Web Cloud
Completed
We are going to modify the increase in IP collection traffic on the Paca region.

We will be taking action on the night of Wednesday 29th to Thursday 30th May from 00:00, and there will be no impact for clients. Sessions terminating in Marseille will terminate in Roubaix during this time.

Update(s):

Date: 2013-05-29 23:12:28 UTC
All ok, the security is fully performing on the LNS.

Date: 2013-05-29 23:08:41 UTC
The routing of the ADVA is ok. We are cabling it all to the LNS in Marseilles.


Date: 2013-05-29 23:07:35 UTC
We are starting the manipulation on the optical buckle.
Posted May 29, 2013 - 09:15 UTC